# Wealth Thermalization Hypothesis

### v2

@@ Abstract
Figure S18: Rescaled reconstructed spectrum $E_{m}/B$ for the data of Hong Kong 2025 (red data points) versus rescaled level number $x=m/N$. The blue curve shows the curve $E_{m}/B=C\,(e^{ax}-1)$ with $C=0.00160\pm 0.00004$ and $a=6.83\pm 0.03$ obtained from a fit in the interval $x\in[0,0.9]$. The value of $a=6.83$ is used in Fig. S17 for the blue curve of the RJE model.
Globally, the fit $E_{m}=C(e^{a(m/N)}-1)$ works rather well at least for some reasonable subinterval. For each data set, we performe two fits of this function for the intervals $x\in[0,0.7]$ and $x\in[0,0.9]$ that provides two interesting values of $a$. We inject these two values of the parameter $a$ in the RJE model and determined which $a$ value gives a better agreement for the Lorenz curve (the value of $\varepsilon$ is determined as usual by matching the Gini coefficient to be identical between initial Lorenz curve and the model curve). Figs. S13S17 show the resulting RJE curves for 5 of our data sets, already discussed in the previous section, with a very good agreement of the RJE model at the optimal fitted values of $a$.
In Fig. S18, we show as illustration one example of an reconstructed spectrum for the data set of the stock market Hong Kong 2025 using the value $N=2683$ and the piecewise linear derivative for $w^{\prime}(h)$. The number $N=2683$ represents the number of companies used in the Hong Kong SE data but the precise choice of this value is not very important and the reconstruction procedure works also nicely for $N=1000$ or $N=10000$ for this example. In this case, the shown fit $E_{m}/B=C\,(e^{ax}-1)$ works quite well for the larger interval $x\in[0,0.9]$ and the resulting value of $a=6.83$ provides a nearly perfect Lorenz curve of the RJE model.
We mention that the bandwidth $B$ of the reconstructed spectrum shown in Fig. S18 is $B=566.6$ and it corresponds to the initial choice $E=1$ to fix the global energy scale such that the rescaled energy of the reconstructed spectrum is $\varepsilon_{rc}=1/B\approx 0.001765$. The value of $\varepsilon_{\rm RJE}$ for the fitted blue curve is slightly modified due to a modified bandwidth of the latter: $\varepsilon_{\rm RJE}\approx\varepsilon_{rc}/[C\,(e^{a}-1)]=0.001193$ which compares to the value $\varepsilon=0.0008381$ given in the caption of Fig. S17 obtained by matching the Gini coefficient. The slightly different value for $\varepsilon$ is due to the matching of the Gini coefficient and the fact that the fit is far from perfect. Furthermore, also the data points for $x>0.9$ are not very reliable.

### v3

@@ B.1 Netscience network
[…]
The modes with limited support length $l(m)\leq 48$ have energies $E_{m}=p/q$ with nice rational values (maximal $q=420$ and other $q\leq 12$). We mention a few examples, such as the mode $E_{25}=-25/12$ with $l(25)=\xi_{\rm IPR}=2$ localized on the two nodes Kim, D. and Goh and with values $\phi^{(25)}_{n}=\pm 1/\sqrt{2}$ for these two nodes. Another example is the pair of two modes $E_{66}=E_{67}=-389/420$ with $l(66)=l(67)=3$ and $\xi_{\rm IPR}=2$ for both. These two modes are localized on the three nodes Rajagopalan, Raghavan, Tomkins.
Figure B.2: $\xi_{\rm IPR}(E_{m})$ of eigenvector $\phi^{(m)}$ of $H$ for the case of the netscience network with $\kappa=0$ and $\kappa=0.5$ versus index $m$ (energy $E_{m}$) in left (right) panel.
Fig. B.2 shows the IPR of eigenvectors for the netscience network at $\kappa=0$ and $\kappa=0.5$ versus mode index $m$ and also mode energy $E_{m}$. The IPR values at $\kappa=0.5$ are roughly a factor 10 larger than the values at $\kappa=0$ and the maximal value of the IPR is 45.98 (94.94) for $\kappa=0$ ($\kappa=0.5$). However, the boundary modes have essentially the same (small) IPR values since due to the large energy gaps these modes are in the quantum perturbative regime even at $\kappa=0.5$ with coupling matrix elements $\sim\kappa/\sqrt{N}=0.5/\sqrt{379}\approx 1/40$ which is much smaller than the boundary energy gaps.
Figure B.3: Lyapunov exponent $\lambda_{m}$ dependence on $E_{m}$ with $m$ being the index of the initial state for the netscience network with $\kappa=0.5$ and $N=379$. $\lambda_{m}$ has been determined from the fit $\ln\|\Delta\psi(t)\|=a+b\ln(t)+\lambda_{m}\,t$. Shown are data from 4 data sets: with 32 selected modes at $t=2^{22}$ for $\beta=0.05$ (red ++ symbol), $\beta=0.2$ (green $\times$ symbol), $\beta=10$ (blue * symbol) and also with 16 selected modes at $t=2^{24}$ for $\beta=10$ (pink $\square$ symbol) obtained from a different computation.
Fig. B.3 shows the Lyapunov exponent $\lambda_{m}$ for selected initial modes $m$ of the netscience network at $\kappa=0.5$ and $\beta=0.05,0.2,10$. The quantity $\|\Delta\psi(t)\|$ mentionned in the figure caption is the vector norm of the difference $\Delta\psi=\psi_{2}(t)-\psi_{1}(t)$ where $\psi_{1}(t)$ and $\psi_{2}(t)$ are both solutions of (II.2) with different but close initial conditions at the initial mode $\psi_{1}(t=0)\approx\psi_{2}(t=0)\approx\phi^{(m)}$. In this case the Lyapunov exponent is defined by $\lambda_{m}=\lim_{t\to\infty}(\ln\|\Delta\psi(t)\|)/t$ and the data in Fig. B.3 has been extracted by the fit $\ln\|\Delta\psi(t)\|=a+b\ln(t)+\lambda_{m}\,t$.
The modes at $\beta=10$ are clearly in the chaotic regime while boundary modes at $\beta=0.2$ and (most) modes at $\beta=0.05$ are in the KAM regime with much smaller (numerical) Lyapunov exponent (we expect that for these states Lyapunov exponents become zero in the limit $t\rightarrow\infty$).
Figure B.4: Color plot of the coarse-grained thermalized occupation probabilities $\rho_{m}=T/(E_{m}-\mu)=(E-\mu)/[N(E_{m}-\mu)]$ for the netscience network (politician network) at $\kappa=0.5$ in left (right) panel. The $x$-axis corresponds to the fraction $m/N\in[0,1]$ with $m=1,\ldots,N$ being the index of energies $E_{m}$ (left to right) and the $y$-axis corresponds to the rescaled energy $\varepsilon=(E-E_{1})/(E_{N}-E_{1})$ (top to bottom for increasing values). The tics indicate integer multiples of 0.1 for both quantities. The color values shown in the color bar correspond to the value of $\rho_{m}$ averaged over intervals of size $1/20$ (for $m/N$ on the $x$-axis) and computed for 21 values $\varepsilon=i/20$, $i=0,1,\ldots,20$ (for the $y$-axis; the minimal value $\varepsilon=0$ has been slightly enhanced and the maximal value $\varepsilon=1$ has been slightly reduced to have a stable computation of the thermalized $\mu$-value). To increase visibility of small values a non-linear color bar scale has been chosen (e.g. green color corresponds to $1/16$).
Fig. B.4 presents coarse-grained color plots in the $m$-$E$ plane of the theoretical RJ-values $\rho_{m}=T/(E_{m}-\mu)$ where $T$ and $\mu$ are determined from the implicit equations (II.16). The horizontal $x$-axis corresponds to $m/N\in]0,1]$ and the vertical $y$-axis to the rescaled energy $\varepsilon=(E-E_{1})/(E_{N}-E_{1})\in[0,1]$. The left (right) panel corresponds to the energy spectrum of $H$ at $\kappa=0.5$ for the netscience network with $N=379$ (politician network with $N=5908$). One clearly sees a condensation on the modes with minimal $m\sim 1$ (maximal $m\approx N$) at $E\approx E_{1}$ ($E\approx E_{N}$) while for intermediate energies the distribution of $\rho_{m}$ is more uniform in $m$.
Figure B.5: The left (right) panel shows the temperature $T$ (the chemical potential $\mu$) versus the energy $E$ for the netscience network at $\kappa=0.5$. The dashed black lines in the right panel correspond to the values of $E_{1}=-6.3829$ and $E_{N}=9.7256$ showing that either $\mu<E_{1}$ (for $T>0$) or $\mu>E_{N}$ (for $T<0$).
Fig. B.5 shows the dependence of $T$ and $\mu$ on $E$ obtained by solving the implicit equations (II.16) with $\rho_{m}=T/(E_{m}-\mu)$. Both curves show the usual behavior with $T>0$ ($T<0$) and $\mu<E_{1}$ ($\mu>E_{N}$) for $E<0$ ($E>0$) with $T\to 0$ for $E\to E_{1}$ or $E\to E_{N}$. Furthermore, for $E\to 0$ we have $|T|\to\infty$ and $|\mu|\sim NT\to\infty$ corresponding to uniform $\rho_{m}\to 1/N$.
It is not very difficult to show that in the limit of large $|\mu|$ the chemical potential is given by the equation
$\mu$ $\approx E+\frac{\sum_{m}(E-E_{m})^{2}}{\sum_{m}(E-E_{m})}\ .$ (B.1)
Here the denominator has the same sign as $E-E_{c}$ where $E_{c}=(\sum_{m}E_{m})/N=\mbox{Tr}(H)/N\sim\kappa/N\approx 0$ is the critical energy (“energy center of mass”) at which the chemical potential and the temperature switch their sign. Here the trace of $H$ is entirely given by the trace of the GOE perturbation $\kappa H_{\rm GOE}$ since Tr(A)=0(A)=0 (the absence of self links implies $A_{nn}=0$).
Note that Figs. B.4 and B.5, which simply provide a generic illustration of the basic properties of RJ-thermalization, are rather similar to Figs. A.2 and A.1 of Appendix A obtained from a different simple uniform spectrum $E_{m}=m/N$.
Figure B.6: Entropy $S_{q}$ versus energy $E$ for the netscience network with $\kappa=0$, $\beta=10$, $N=379$. Shown are 64 selected modes at $t=2^{24}$ (red ++ symbols) and $t=2^{20}$ (green $\times$ symbols). The blue line shows the energy dependence of the theoretial thermalized entropy. $S_{q}$ has been computed by Equation (II.17) using $\rho_{m}$ values obtained as the time average $\rho_{m}=\langle|C_{m}(\tilde{t})|^{2}\rangle$ for $t/2<\tilde{t}\leq t$ (for $t=2^{24}$ or $t=2^{20}$ according to the selected data in this figure).
For the case of the netscience network with $\kappa=0$ and $\beta=10$, we show in Fig. B.6 the energy dependence of $S_{q}$. Here the agreement with the theoretical curve is similar as for the case with $\kappa=0.5$ but for this longer iteration time scales $t=2^{24}$ are required.
Figure B.7: As Figure II.3 for the netscience network with $\kappa=0.5$, $\beta=4$, $N=379$ and the two initial modes $m_{0}=8,54$. The blue curve shows the RJ theoretical curve $\rho_{\rm RJ}(E)=T/(E-\mu)$. The values of $T$, $\mu$ and $\langle E\rangle$ for the 2 initial modes $m_{0}=8,54$ are $T=0.008006,0.01461$, $\mu=-6.4,-6.574$ and $\langle E\rangle=-3.366,-1.038$.
In Fig. B.7, we present two examples of nicely thermalized states for the netscience network with $\kappa=0.5$ and $\beta=4$, Due to the smaller coupling value of $\beta$ the fluctuations at the shorther time $t=2^{22}$ are stronger than for similar states at $\beta=10$.
Figure B.8: Results for the netscience network with $\kappa=0.5$ and two initial states localized either on the node of Barbasi or Newman. Top panels show the time dependence of $S_{q}$ and $S_{B}/N$ for $\beta=4$ (left) and $\beta=10$ (right). The two top (bottom) curves correspond to $S_{B}(t)/N$ ($S_{q}(t)$) and dashed lines indicate the theoretical thermalized values for each case. The 6 bottom panels show $\rho_{m}$ versus mode energy $E_{m}$ for both initial conditions and $\beta=10,4$ ($t=2^{22},2^{24}$) and $\beta=1$ ($t=2^{23},2^{25}$). See text and captions of Figs. II.3-II.5 for more technical details.
Fig. B.8 shows some results for the intial condition localized on the nodes of Barabasi or Newman (top two PageRank nodes for the netscience network). In its first two top panels we see the time dependence of $S_{q}(t)$ and $S_{B}(t)/N$ for both cases and the two values $\beta=4$ and $\beta=10$. All eight curves saturate to the corresponding theoretical thermalized entropy values which are obtained from the usual value for $E=\langle E\rangle=\sum_{m}E_{m}\rho_{m}(t=2^{24})$ to determine $T$ and $\mu$ and with the numerical final values of $\rho_{m}(t=2^{24})$. For the case $\beta=4$ the onset of thermalization is somewhat delayed in the initial phase with smaller entropy values for $t<2^{14}$ (in comparison to the case $\beta=10$). For such states the conserved energy in (II.3) is given by the simple formula $E=\mathcal{H}=H_{n_{0},n_{0}}+\beta/2\approx\beta/2$ since $H_{n_{0},n_{0}}\sim 1/\sqrt{N}$ (absence of self links with $A_{n_{0},n_{0}}=0$ and small diagonal matrix element from the GOE perturbation).
The 6 panels in the bottom three rows of Fig. B.8 show, in the same style as Fig. II.3, for different states the dependence of $\rho_{m}$ on $E_{m}$ for the largest iterations times $t=2^{24}$ ($\beta=10,4$) or $t=2^{25}$ ($\beta=1$) (red ++ symbols) and at a shorter time $t/100$ (green $\times$ symbols). Furthermore, each panel also presents the theoretical curve $\rho_{\rm RJ}(E_{m})=T/(E_{m}-\mu)$ (blue line) where $T$ and $\mu$ are obtained by solving the constraints (II.16) at $E=\langle E\rangle=\sum_{m}E_{m}\rho_{m}(t)$ for the numerical data of $\rho_{m}(t)$ at the largest available value of $t$. The obtained values of $T$, $\mu$ and $\langle E\rangle$ for the 3 cases $\beta=10,4,1$ of Barabasi are T=0.01322,0.02075,0.02822T=-0.01322,-0.02075,-0.02822, $\mu=9.754,9.853,11.2$ and $\langle E\rangle=4.744,1.99,0.5012$. For Newman these values (at same $\beta$-order) are T=0.01292,0.02074,0.02811T=-0.01292,-0.02074,-0.02811, $\mu=9.753,9.853,11.16$ and $\langle E\rangle=4.855,1.991,0.5069$. The negative values of $T$ are coherent with the positive energy values above the critical value $E_{c}\approx 0$ for the transition from positive to negative $T$ (see also Eq. (B.1) and the discussion of Fig. B.4 above).
From the physical point of view, we see that globally for all cases the states thermalize well to the theoretical curve for (nearly) all values $\rho_{m}$. The data points for the shorter time value $t/100$ show stronger fluctuations as expected and for $\beta=1$ the quality of convergence is also a bit reduced, probably larger iterations times are still needed here.
B.2 Politician network
Figure B.9: $\xi_{\rm IPR}(E_{m})$ of eigenvector $\phi^{(m)}$ of $H$ for the case of the politician network with $\kappa=0$ and $\kappa=0.5$ versus index $m$ (energy $E_{m}$) in left (right) panel.
Fig. B.9, shows the IPR values of eigenmodes for the politician network at $\kappa=0$ and $\kappa=0.5$ (in the same style as Fig. B.2). For a significant number of boundary modes the IPR is very small with sames values between $\kappa=0$ and $\kappa=0.5$. Due to the large boundary energy gaps the states are essentially the same for both $\kappa$ values. Modes in the bulk have a quite large distribution of IPR values between localized states (IPR $\sim 10$) and maximal values IRP $\sim 1200$ (for both $\kappa$ values). There is also a signifiant reduction of typical IPR values for energies $E_{m}\approx 0$ by a factor 3 (2) for $\kappa=0$ ($\kappa=0.5$). Furthermore, the probability (density of data points) to find very small IPR values (for center modes) is strongly reduced for $\kappa=0.5$ in comparison to $\kappa=0$.
Figure B.10: Entropy $S_{q}$ ($S_{B}/N$) versus energy $E$ in left (right) panel for some initial modes of the politician network with $\kappa=0.5$, $N=5908$, $\beta=10$. For each case three data points for the three longest time values are shown to illustrate the thermalization. The first four modes ($m_{0}=8,16,32,128$) were computed with the time step $dt=1/32$ for the symplectic integrator up to $t=2^{19}$. For other modes $S_{q}$ and $S_{B}$ were computed with $dt=1/16$ (for $128<m_{0}\leq 600$) or $dt=0.1$ (for $600<m_{0}$) both up to $t=2^{20}$. The blue line shows the energy dependence of the theoretical thermalized entropy for both entropy quantities.
In Fig. B.10, we show for a selected number of modes with initial energies in the interval $14<E_{m_{0}}<0$ the dependence of both entropy values $S_{q}$ and $S_{B}/N$ on the final (linear) energy $E=\langle E\rangle=\sum_{m}E_{m}\rho_{m}$ (which is typically very close to $E_{m_{0}}$). For each data point the values for three successive time values $t=2^{l-2},2^{l-1},2^{l}$ (with either $l=19$ for $m_{0}\leq 128$ or $l=20$ for $m_{0}>128$) are shown indicating a clear tendency for convergence to the theoretical curve. For three of the four modes with $m_{0}\leq 128$ and $S_{q}$ the last data points are actually already very close to the theoretical curve and the remaining mode at $E\approx-12$ shows a strong increase of $S_{q}$ between the last two time values indicating a potential convergence at later times.
B.3 Entropy in the RMT model
Figure B.11: Time dependence of $S_{q}$ for the case of a GOE matrix ($N=64$, semicircle radius 1) using the data of [17]. The full lines correspond to $S_{q}(t)$ for the cases of Figure S1 of [17] and the dotted lines indicate the theoretical thermalized RJ values (using the color for the cases with $\beta=1$). This figures is very similar to Figure S1 of [17]. It is shown here for convenience.
Figure B.12: Time dependence of $S_{B}/N$ for the case of a GOE matrix ($N=64$, semicircle radius 1) using the data of [17]. The full lines correspond to $S_{B}(t)/N$ for the cases of Figure S1 of [17] and the dotted lines indicate the theoretical thermalized RJ values (using the color for the cases with $\beta=1$). $S_{B}$ has been computed by (II.20) using $h_{B}=1/N^{2}$.
Figs. B.11 and B.12 show the time dependence of $S_{q}(t)$ and $S_{B}(t)/N$ respectively for a pure GOE matrix (corresponding to $A=1$ and $\kappa=1$ for $N=64$) for the cases already shown in Fig. S1 of [17] which is actually very similar to Fig. B.11 (the latter shows more theoretical values as well). For $S_{q}$ we have a nonmonotonic time dependence for the initial states at $m_{0}=2,3$ at intermediate times. However, at large times these modes are well thermalized and close to the RJ condensate phase. For $S_{B}$ the time dependence is always monotonic in the chaotic regime.
B.4 Wealth inequality and Lorenz curves
Figure B.13: Density of states (Lorenz curves) for the netscience network at $\kappa=6$ in left (right) panel.
Fig. B.13 shows the density of state for the netscience network and a set of Lorenz curves for the case of a very strong value of $\kappa$. Now, the density of states is rather close to the semicercle law with radius $\kappa=6$ and the behavior for $h>h_{0}$ ($h_{0}=$ maximal value at which $w(h)=0$) corresponds more to a smooth curve instead of a straight line as for the case $\kappa=0.5$ shown Fig. II.8.